what is the maximum speed at which a car can negotiate an unbanked turn (radius=60.0 m) in dry weather (coefficient of static friction = 0.900)?
Answer #1

Using Force balance on Car at unbanked turn

F_c = F_f

Centripetal force due to turn = Friction force

m*V^2/R = *N

N = Normal force on car = m*g

m*V^2/R = *m*g

V_max = sqrt (*g*R)

V_max = sqrt (0.900*9.81*60.0)

V_max = 23.0 m/sec = max speed at which car can turn in dry weather

Please Upvote.